The scores on the entrance exam at a well-known, exclusive law school are normally distributed with a mean score of 250 and a standard deviation equal to 89. At what value should the lowest passing score be set if the school wishes only 2.5 percent of those taking the test to pass? (Round your answer to nearest whole number.)
Answer #1

Using standard normal table,

P(Z < z) =2.5 %

= P(Z < z) = 0.025  

= P(Z <-1.96 ) = 0.025

z = -1.96 Using standard normal z table,

Using z-score formula  

x= z * +

x= -1.96 *89+250

x= 75.56

x=76